OpenStudy (anonymous):

A fraction reduces to 36. If its denominator is 6x5, what is its numerator?

OpenStudy (cwrw238):

just multiply by 36

OpenStudy (anonymous):

I meant to say that it's denominator is 6x^5

OpenStudy (cwrw238):

right - i assumed that to be the case all you nee to do is calculate 36 * 6x^5

OpenStudy (anonymous):

If I give you the answer choices, can you help to explain further to me about which one is right? Because I tried the lesson, and I was extremely confused

OpenStudy (cwrw238):

ok

OpenStudy (anonymous):

63x 63x5 6x5 67x5

OpenStudy (anonymous):

Sorry it took so long, my internet was acting up...

OpenStudy (cwrw238):

i'm confused - the answer is 216 x^5

OpenStudy (cwrw238):

ohh! unless its (b) which should be written 6^3x^5

OpenStudy (cwrw238):

because 6^3 = 36*6 = 216

OpenStudy (anonymous):

Ohhh! Okay, that makes a lot of sense. Thank you so much!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

While you are here, could you help me with one more?(:

OpenStudy (anonymous):

Do you have the answers to this test? I want to check mine. I cannot figure out this question